not sure of the appropriate topic for shifting:
given f(x)=x^2-8
shift f(x)2 units to the right I got h(x)=f(x-c)=(x-2)^2-8
shift f(x)2 units down I got h(x)=x^2+(-8-2)=x^2-6
shift f(x)2 units right and 2 units down I got h(x)=(x-2)^2+(-8-2)
Am I correct?

Answer by ice_water35(8)

>>>>>This is correct:
shift f(x)2 units to the right I got h(x)=f(x-c)=(x-2)^2-8
shift f(x)2 units down I got h(x)=x^2+(-8-2)=x^2-6
-in this case, your solution is correct but your final answer is wrong because -8 - 2 is -10. So, your solution and answer must be like this:
I got h(x)=x^2+(-8-2)=x^2-10
shift f(x)2 units right and 2 units down I got h(x)=(x-2)^2+(-8-2)
-in this case, your answer is not the final answer. we may simplify this, so that your answer must be like this:
I got h(x)=(x-2)^2-10
I change some signs in your questions.
If you have further questions, just contact me. This is my number: 09223011043